If perimeter of pool is 576 ft. The width is 1/3 its length, what are the dimensions?

Answer by macston(5194)

.
L=length; W=width=(1/3)L; P=perimeter=576ft
P=2(L+W) Substitute for W.
576ft=2(L+(1/3)L) Divide each side by 2.
288ft=(4/3)L
288ft=(4/3)L Multiply each side by 3/4.
216ft=L ANSWER 1: The length is 216 feet.
W=(1/3)L=(1/3)216ft=72 ft ANSWER 2: The width is 72 feet.
CHECK:
P=2(L+W)
576ft=2(216ft+72ft)
576ft=2(288ft)
576ft=576ft